# WHALE-FL: Wireless and Heterogeneity Aware Latency Efficient Federated Learning over Mobile Devices via Adaptive Subnetwork Scheduling

### v1

@@ 4.2 Utility Value to Subnetwork Size Conversion
Algorithm 1 WHALE-FL Procedure
1:Input: Data $\{D_{1},\cdots,D_{M}\}$ distributed on $M$ mobile devices, the number of local epochs $E$, the local minibatch size $B$, the learning rate $\eta$, the channel shrinkage ratio $s$, and the number of subnetwork size levels $P$.
2:Initialize global model $W(g,0)$
3:for each communication round $r=0,1,2,\cdots$ do
4:     for each client $i\in[M]$ in parallel do
5:         $s_{i}\leftarrow$ SubnetSelect$(SE_{i,r},TE_{i,r},W_{i}^{max},P)$
6:         $d_{i}\leftarrow s_{i}d_{g},k_{i}\leftarrow s_{i}k_{g}$
7:         $W(i,r)\leftarrow W(g,r)[:d_{i},:k_{i}]$
8:         $W(i,r+1)\leftarrow$ ClientUpdate$(i,s_{i},W(i,r))$
9:     end for
10:     for each size level $p\in[P]$ do
11:         Compute $W_{g}^{p,r+1}$ via Eqn. (2.2)
12:     end for
13:     Update the global model $W(g,r+1)$ via Eqn. (9)
14:end for
15:SubnetSelect$(SE_{i,r},TE_{i,r},W_{i}^{max},P)$:
16:if $r=0$ then
17:     $s_{i}\leftarrow 1$
18:else
19:     $Util(i,r)\leftarrow SE_{i,r}\times\left(TE_{i,r}\right)^{\beta}$
20:     Compute $U_{n}(i,r)$ via Eqn. (6)
21:     Compute $\hat{W}(i,r)$ via Eqn. (8)
22:     Compute $W(i,r)$ via Eqn. (7) and determine the corresponding $p$
23:     $s_{i}\leftarrow s^{p-1}$
24:end if
25:Return $s_{i}$
26:ClientUpdate$(i,s_{i},W(i,r))$:
27:$B_{i}\leftarrow$ Split local data $D_{i}$ into batches of size $B$
28:for each local epoch $e$ from $1$ to $E$ do
29:     for batch $b_{i}\in B_{i}$ do
30:         $W(i,r+1)\leftarrow W(i,r)-\eta\nabla l(W(i,r),s_{i};b_{i})$
31:     end for
32:end for
33:Return $W(i,r+1)$
Then, mobile devices conduct local computing according to their selected subnetworks, respectively, followed by transmitting local model updates to FL server. Following the same aggregation method in [9], FL server aggregates updated local models with heterogeneous subnetworks and updates the global model as
@@
In summary, during FL training, mobile devices collect their local information at runtime, including up-link channel quality, background computational loads, memory usage, training loss, etc. Based on the collected information, at the beginning of the $r$-th training round, each device leverages Eqn. (5) to trade-off system efficiency and training efficiency, and calculates its adaptive subnetwork selection utility value $Util(i,r)$. The utility value is then normalized into $U_{n}(i,r)$. Device $i$ uses $U_{n}(i,r)$ to determine the subnetwork size and select an appropriate subnetwork for its local training according to Eqn. (7) and Eqn. (8). After that, FL server aggregates locally trained subnetworks with different sizes and updates the global model for the next round training. The pseudocode of WHALE-FL procedure is provided in Alg. 1.
5 Experimental Setup
5.1 WHALE-FL Testbed
@@
The WHALE-FL testbed consists of an FL aggregator and a set of heterogeneous mobile devices as FL clients. A NVIDIA RTX 3090 serves as the FL server, whose memory capacity is $24$ GB. For heterogeneous FL clients, we have incorporated $5$ types of mobile devices, i.e., MacBookPro2018, NVIDIA Jetson Xavier, NVIDIA Jetson TX2, NVDIA Jetson Nano, and Raspberry Pi 4, representing a range of on-device computing capabilities from high to low. The WHALE-FL system involves a total of $20$ mobile devices, $4$ devices per type. Communication between FL clients and the FL server is facilitated through LTE, BlueTooth, and Wi-Fi $5$ transmission environments. The corresponding transmission rates are $80$ Mbps (Wi-Fi $5$), $20$ Mbps (LTE), and $10$ Mbps (BlueTooth $3.0$), respectively. We set hidden channel shrinkage ratio $s=\frac{1}{2}$ and adopt $5$ subnetwork size levels. Accordingly, the model shrinkage ratios for the $5$ size levels (i.e., $p=1,2,\cdots,5$) are $1$, $\frac{1}{4}$, $\frac{1}{16}$, $\frac{1}{64}$, and $\frac{1}{256}$, respectively.
5.2 Datasets, Models, Parameters and Baselines
We conduct our experiments with three different FL tasks: image classification, human activity recognition and language modeling. As for the image classification task, we train a CNN on MNIST dataset [7] and a ResNet18181818 on CIFAR10101010 dataset [12]. Human activity recognition involves training a CNN on the HAR dataset [10], and a Transformer is trained on the WikiText2 dataset [8] for the language modeling task. We use the balanced non-IID data partition [16]. Take the MNIST dataset as example, the total number of classes is $10$. Our default setup is that each device has $\sigma=2$ classes. We apply similar non-IID setup to other tasks. Besides, we set $\Delta_{th}$ to follow the typical learning rate. As for $l_{target}$, it is the only task-specific hyperparameter. We let $l_{target}$ be $0.001$ for CNN@MNIST and CNN@HAR, $0.1$ for Resnet18181818@CIFAR10101010 and $1$ for Transformer@WikiText2222 in our experiments.
We compare our WHALE-FL against two peer designs across these FL tasks: i) FedAvg [20], where all the clients train with full-sized models, and ii) HeteroFL [9], where subnetwork assignments are fixed and align with clients’ full computation and communication capabilities.
@@ 5.2 Datasets, Models, Parameters and Baselines
6 Evaluation and Analysis
6.1 Latency Efficiency and Learning Performance
@@
As the results shown in Fig. 2, the proposed WHALE-FL consistently achieves remarkable training speedup across various FL tasks without sacrificing learning accuracy. Compared with FedAvg, WHALE-FL accelerates the FL training to the target testing accuracy by approximately 1.5x, 1.9x, 1.3x and 2.1x for FL tasks including CNN@MNIST, ResNet18181818@CIFAR10101010, Transformer@WikiText2, and CNN@HAR, respectively. As detailed in Sec. 3, HeteroFL’s static fixed-size subnetwork assignment policy is not aware of system and training dynamics, which may slow down FL convergence. In contrast, considering both system efficiency and training efficiency, WHALE-FL appropriately assesses the subnetwork selection utility for individual device and adaptively adjusts the local subnetwork size to suit for time-varying communication and computational conditions and dynamic changing requirements of FL training at different FL training stages, in order to reduce training latency. Consequently, compared with HeteroFL, WHALE-FL achieves a notable speedup of 1.2x, 1.3x, 1.2x and 1.5x for the tested 4 learning tasks, respectively.
Figure 3: System efficiency utility and subnetwork size level changes over training time (Macbookpro2018, CNN@HAR).
@@
Figure 4: Performance comparison: WHALE-FL, system efficiency only and training efficiency only designs (ResNet18181818@CIFAR10101010).
6.2 Subnetwork Size Changes over Rounds
As shown in Fig. 3, the subnetwork size for Macbookpro2018 (CNN@HAR) dynamically changes during FL training process. At the early training stage, the subnetwork size changes along with system efficiency utility. At the middle training stage, the subnetwork size gradually increases jointly determined by system and training efficiency. At the late stage, the subnetwork size reduces for better training latency efficiency. Here, we take Macbookpro2018 for example, and the analysis applies to all participating mobile devices.
6.3 System Efficiency vs Training Efficiency
@@
To differentiate system efficiency’s contributions from training efficiency’s ones, we compare WHALE-FL with system efficiency utility only and training efficiency utility only schedulings. As the results shown in Fig. 4, WHALE-FL converges faster than training efficiency only subnetwork scheduling when achieving the target accuracy, since training efficiency only design has no consideration of system dynamics and its impacts on subnetwork size selection; WHALE-FL has better testing accuracy but proceeds slower than system efficiency only subnetwork scheduling at the early training stage. The reason behind is that system efficiency only design prioritizes system dynamics while ignoring dynamic model accuracy requirements for local training at different FL training stages. WHALE-FL trades-off system and training efficiencies and jointly considers their benefits for FL training.
Figure 5: Sensitivity performance under different $\alpha$, $\beta$, and $\gamma$ value configurations (CNN@MNIST).
@@
Figure 6: Sensitivity performance under different $\alpha$, $\beta$, and $\gamma$ value configurations (Transformer@WikiText2).
6.4 Sensitivity Analysis
We further evaluate the impacts of $\alpha$, $\beta$ and $\gamma$, defined in the subnetwork selection utility function, on subnetwork scheduling. Taking CNN@MNIST as an example, we conduct sensitivity analysis of its performance under different $\alpha$, $\beta$ and $\gamma$ values, and present the results in Fig. 5. For generalization purpose, we have also conducted the sensitivity study for NLP task, whose results are shown in Fig. 6.
Figure 7: Sensitivity performance under different $U_{th}$ value configurations (Transformer@WikiText2222).
The $\alpha$-value helps to separate the middle and late FL training stages, which determines the increment/decrement of selected subnetwork sizes. If $\alpha=1$, FL training stops when ($l_{i,r-1}\leq l_{target}$) as shown in Eqn. (5) and there will be no late stage and corresponding subnetwork size decrease. Thus, as shown in Fig. 5, compared with $\alpha=2$ or $5$ cases, $\alpha=1$ based FL training is a little bit slow to converge.
@@ 6.4 Sensitivity Analysis
The hyperparameter $\beta$ trades-off system efficiency and training efficiency utilities. The large/small $\beta$ value means that the device prioritizes training/system efficiency. As the results shown in Fig. 5, we find that the FL training converges slower but achieves higher testing accuracy when $\beta$ is large, e.g., $\beta=5$ in Fig. 5, while FL training is faster at early stage but achieves lower testing accuracy when $\beta$ is smaller, e.g., $\beta=1$ in Fig. 5. System efficiency and training efficiency are somehow balanced when $\beta=2$ in Fig. 5. Thus, although $\beta$ is a developer-specified factor, a proper selection of $\beta$ value helps FL training converge fast while achieving good learning performance.
Figure 5 reveals the training performance is not that sensitive to $\gamma$ values. Recall that $\gamma$ is the step-size for increment/decrement of the training efficiency utility. The results reflect that our WHALE-FL consistently outperforms the baseline, FedAvg, regardless of how the increment/decrement step-size is configured in Eqn. (5). This may be attributed to the proposed normalization and discretization operations on utility values in Sec. 4.2, which mitigates the impacts of $\gamma$ on the latency or learning performance. For developer-specified $U_{th}$, it can serve as a cross-device guidance parameter for selecting subnetwork sizes. The impact of $U_{th}$ is shown in Fig. 7. Increasing $U_{th}$ makes all clients opt for smaller subnetworks, and decreasing $U_{th}$ makes all clients tend to select bigger subnetworks.
6.5 Impacts of Data Heterogeneity
We further evaluate the impacts of data heterogeneity on WHALE-FL’s performance. Here, we take CNN@MNIST as an example and use the balanced non-IID data partition [16]. The total number of classes in the MNIST dataset is $10$. We study the cases that each device has $\sigma=2,5$ or $10$ classes, where the data distribution is IID if $\sigma=10$, i.e., every device has all classes. The results are shown in Table 1, where we find that (i) FL training with non-IID data takes longer time to converge, and (ii) embracing both system and training efficiency utilities, WHALE-FL can remarkably improve FL training delay efficiency compared with HeteroFL and FedAvg under various data heterogeneity scenarios.
Table 1: Performance comparison under different data heterogeneity levels (CNN@MNIST), where “OL" represents the overall latency.
@@ 6.5 Impacts of Data Heterogeneity
Target Acc. 90% 95% 97%
Metric Hours (OL) Hours (OL) Hours (OL)
WHALE 0.82 (1.5x) 0.24 (1.4x) 0.08 (3.8x)
HeteroFL 0.98 (1.3x) 0.26 (1.3x) 0.10 (3.0x)
FedAvg 1.23 (1.0x) 0.33 (1.0x) 0.30 (1.0x)
7 Conclusion

### v3

@@ Introduction
Then, mobile devices conduct local computing according to their selected subnetworks, respectively, followed by transmitting local model updates to the FL server. Following the same aggregation method in (Diao, Ding, and Tarokh 2021), the FL server aggregates updated local models with heterogeneous subnetworks and updates the global model as
@@
In summary, during FL training, mobile devices collect their local information at runtime, including up-link channel quality, background computational loads, memory usage, training loss, etc. Based on the collected information, at the beginning of the $r$-th training round, each device leverages Eqn. (Adaptive Subnetwork Selection Utility) to trade-off system efficiency and training efficiency, and calculates its adaptive subnetwork selection utility value $Util(i,r)$. The utility value is then normalized into $U_{n}(i,r)$. Device $i$ uses $U_{n}(i,r)$ to determine the subnetwork size and select an appropriate subnetwork for its local training according to Eqn. (8) and Eqn. (9). After that, FL server aggregates locally trained subnetworks with different sizes and updates the global model for the next round training. In the appendix, we provide the convergence analysis for WHALE-FL based on (Wang et al. 2023), and show that WHALE-FL can converge under adaptive subnetwork size scheduling.
Experimental Setup
WHALE-FL Testbed
@@
The testbed consists of an FL aggregator and a set of heterogeneous mobile devices as FL clients. A NVIDIA RTX 3090 serves as the FL server, whose memory capacity is $24$ GB. For heterogeneous FL clients, we have incorporated $5$ types of mobile devices, i.e., MacBookPro2018, NVIDIA Jetson Xavier, NVIDIA Jetson TX2, NVDIA Jetson Nano, and Raspberry Pi 4, representing a range of on-device computing capabilities from high to low. The WHALE-FL system involves a total of $20$ mobile devices, $4$ devices per type. Communication between FL clients and the FL server is facilitated through LTE, BlueTooth, and Wi-Fi $5$ transmission environments. The corresponding transmission rates are $80$ Mbps (Wi-Fi $5$), $20$ Mbps (LTE), and $10$ Mbps (BlueTooth $3.0$), respectively. We set hidden channel shrinkage ratio $s=\frac{1}{2}$ and adopt $5$ subnetwork size levels. Accordingly, the model shrinkage ratios for the $5$ size levels (i.e., $p=1,2,\cdots,5$) are $1$, $\frac{1}{4}$, $\frac{1}{16}$, $\frac{1}{64}$, and $\frac{1}{256}$, respectively.
Datasets, Models, Parameters and Baselines
We conduct our experiments with three different FL tasks: image classification, human activity recognition and language modeling. As for the image classification task, we train a CNN on MNIST dataset (Deng 2012) and a ResNet18181818 on CIFAR10101010 dataset (He et al. 2015). Human activity recognition involves training a CNN on the HAR dataset (Gupta et al. 2022), and a Transformer is trained on the WikiText2 dataset (Devlin et al. 2019) for the language modeling task. We use the balanced non-IID data partition (Li et al. 2021b). Take the MNIST dataset as an example, the total number of classes is $10$. Our default setup is that each device has $\sigma=2$ classes. We apply a similar non-IID setup to other tasks. The Fisher information’s window size $|\mathbf{D}|$ = 10. We employ the following peer designs for performance evaluation: (i) FedAvg (McMahan et al. 2017), where all the clients train with full-sized models; (ii) HeteroFL (Diao, Ding, and Tarokh 2021), where subnetwork assignments are fixed and aligned with clients’ full computation and communication capabilities; (iii) FedDropout (Wen, Jeon, and Huang 2021), which generates subnetworks by choosing the neurons at random; and (iv) FedRolex (Alam et al. 2024), which uses a rolling subnetwork extraction scheme in each FL training round. In particular, we compare the peer design with WHALE-FL’s corresponding extension, i.e., the integration of the peer design and WHALE-FL, e.g., FedRolex vs WHALERolex.
@@
Task CV NLP HAR
CNN@MNIST Resnet@CIFAR10 Transformer@Wikitext2 CNN@HAR
Target Accuracy $85\%$ $70\%$ $37\%$ $88\%$
Method Speedup
WHALE vs HeteroFL 1.74x 1.25x 1.21x 1.06x
WHALERolex vs FedRolex 1.75x 1.32x 1.24x 1.10x
WHALEDropout vs FedDropout 1.70x 1.24x 1.20x 1.05x
Table 1: Performance comparison under different subnetwork methods (Speedup).
Task CV NLP HAR
CNN@MNIST Resnet@CIFAR10 Transformer@Wikitext2 CNN@HAR
Method Final Accuracy Improvement
FedAvg 92.71% 80.61% 40.54% 92.94%
HeteroFL $\Rightarrow$ WHALE 87.42% $\Rightarrow$ 89.29% 71.65% $\Rightarrow$ 75.32% 37.40% $\Rightarrow$ 39.28% 88.86% $\Rightarrow$ 91.38%
FedRolex $\Rightarrow$ WHALERolex 87.82% $\Rightarrow$ 89.87% 72.52% $\Rightarrow$ 79.57% 38.02% $\Rightarrow$ 39.66% 89.11% $\Rightarrow$ 92.03%
FedDropout $\Rightarrow$ WHALERolex 86.16% $\Rightarrow$ 87.52% 70.08% $\Rightarrow$ 73.45% 37.19% $\Rightarrow$ 39.06% 88.25% $\Rightarrow$ 89.55%
Table 2: Performance comparison under different subnetwork methods (Final Accuracy Improvement).
Evaluation and Analysis
Latency Efficiency and Learning Performance
@@
As the results shown in Fig. 1, the proposed WHALE-FL consistently achieves remarkable training speedup across various FL tasks without sacrificing learning accuracy. Compared with FedAvg, WHALE-FL accelerates the FL training to the target testing accuracy by approximately 1.5x, 1.9x, 1.3x, and 2.1x for FL tasks including CNN@MNIST, ResNet18181818@CIFAR10101010, Transformer@WikiText2, and CNN@HAR, respectively. As detailed in Sec. Motivation, HeteroFL’s static fixed-size subnetwork assignment policy is not aware of system and training dynamics, which may slow down FL convergence. In contrast, considering both system efficiency and training efficiency, WHALE-FL appropriately assesses the subnetwork selection utility for individual devices and adaptively adjusts the local subnetwork size to suit for time-varying communication and computational conditions and dynamic changing requirements of FL training at different FL training stages, in order to reduce training latency. Consequently, compared with HeteroFL, WHALE-FL achieves a notable speedup of 1.74x, 1.25x, 1.21x and 1.06x for the tested 4 learning tasks, respectively. Results in Tables 1 and 2 further demonstrate that WHALE-FL and WHALE-FL based extensions (i.e., WHALEDropout and WHALERolex) achieve faster convergence and better testing accuracy than the peer designs across different FL tasks.
Subnetwork Size and Fisher Information Changes
Figure 2: Fisher information and subnetwork size level changes over training time (CNN@MNIST). From left to right, the performance of the user-side models on MacBookPro 2018, NVIDIA Jetson TX2, and Raspberry Pi 4, as well as the global model’s performance, are shown.
As shown in Fig. 2, across the three heterogeneous devices - MacBookPro 2018 (high-end), NVIDIA Jetson TX2 (medium), and Raspberry Pi 4 (low-end) - the subnetwork sizes adjust following the $|\mathbf{D}|$-averaged changes of local Fisher information. The results align with our expectations: When Fisher information is high, the subnetwork size increases to enhance the global model’s accuracy; as training proceeds and Fisher information decreases, indicating that its impacts on learning decrease, the subnetwork is becoming smaller to improve the training time efficiency. On the server side, the averaged size of the aggregated local subnetworks changes along with the global model’s Fisher information, which exhibits a similar trend to the local Fisher information. Figure 2 demonstrates that WHALE-FL effectively captures training dynamics while selecting appropriate subnetwork sizes for heterogeneous devices.
@@
Figure 3: Performance comparison of WHALE-FL, system efficiency only and training efficiency only designs (ResNet18181818@CIFAR10101010).
System Efficiency vs Training Efficiency
@@
To differentiate system efficiency’s contributions from training efficiency’s ones, we compare WHALE-FL with system efficiency utility only and training efficiency utility only schedulings. As the results shown in Fig. 3, WHALE-FL converges faster than training efficiency only subnetwork scheduling when achieving the target accuracy, since training efficiency only design has no consideration of system dynamics and its impacts on subnetwork size selection; WHALE-FL has better testing accuracy but proceeds slower than system efficiency only subnetwork scheduling at the early training stage. The reason behind is that the system efficiency only design prioritizes system dynamics while ignoring dynamic model accuracy requirements for local training at different FL training stages. WHALE-FL trades-off system and training efficiencies and jointly considers their benefits for FL training.
(a) $\beta$, CNN@MNIST.
(b) $D$, CNN@MNIST.
(c) $U_{th}$, CNN@MNIST.
(d) $\beta$, Transf.@WikiText2.
(e) $D$, Transf.@WikiText2.
(f) $U_{th}$, Transf.@WikiText2.
@@
Figure 4: Sensitivity analysis under different $\beta$, $D$, and $U_{th}$ values (a-c: CNN@MNIST; d-e: Transformer@WikiText2).
Sensitivity Analysis
We further evaluate the impacts of $\beta$, $D$ = $|\mathbf{D}|$, and $U_{th}$, defined in the subnetwork selection utility function, on subnetwork scheduling.
@@
The hyperparameter $\beta$ trades-off system efficiency and training efficiency utilities. The large/small $\beta$ value means that the device prioritizes system/training efficiency. As the results shown in Fig. 4(a) and Fig. 4(d), we find that the FL training converges slower but achieves higher testing accuracy when $\beta$ is small, e.g., $\beta=1$, while FL training is faster at early stages but achieves lower testing accuracy when $\beta$ is larger, e.g., $\beta=5$. System efficiency and training efficiency are somehow balanced when $\beta=2$. Thus, although $\beta$ is a developer-specified factor, a proper selection of $\beta$ value helps FL training converge fast while achieving good learning performance. The hyperparameter $D$ represents the window size for calculating the averaged Fisher information. A small window size, such as $D=1$ in Fig. 4(b) and Fig. 4(e), makes the subnetwork size updates sensitive to changes in Fisher information, leading to fluctuations in model accuracy during training. Conversely, employing a larger window size, like $D=20$, results in slower subnetwork-size changes. This may cause a situation where a small-sized subnetwork is well-trained while the clients have no chance to switch to the larger-sized subnetworks, thus impairing the training performance during the critical learning periods. A window size of $D=10$ strikes a good balance, achieving faster convergence. Similarly, a higher $U_{th}$, e.g., $U_{th}$ = 50 shown in Fig. 4(c) and  4(f), leads clients to choose smaller subnetworks, which speeds up FL convergence in the early stages by reducing transmission and computation delays but results in lower final accuracy. Conversely, with $U_{th}$ = 10, clients select larger subnetworks, which slows down convergence but yields higher accuracy. A proper $U_{th}$ selection helps to balance learning performance and delay efficiency.
@@
Target Acc. 85% 90% 95%
Metric Hours (SP)
FedAvg 1.12 (1.00x) 0.33 (1.00x) 0.30 (1.00x)
HeteroFL 1.06 (1.06x) 0.26 (1.27x) 0.10 (3.00x)
WHALE 0.61 (1.84x) 0.19 (1.74x) 0.08 (3.75x)
FedDropout 1.09 (1.03x) 0.28 (1.18x) 0.11 (2.73x)
WHALEDropout 0.64 (1.75x) 0.21 (1.57x) 0.08 (3.75x)
FedRolex 1.03 (1.09x) 0.25 (1.32x) 0.10 (3.00x)
WHALERolex 0.59 (1.90x) 0.18 (1.83x) 0.07 (4.29x)
Table 3: Performance comparison under different data heterogeneity (CNN@MNIST), where “SP” is the speedup.
Impacts of Data Heterogeneity
We further evaluate the impacts of data heterogeneity on WHALE-FL’s performance. Here, we take CNN@MNIST as an example and use the balanced non-IID data partition (Li et al. 2021b). The total number of classes in the MNIST dataset is $10$. We study the cases where each device has $\sigma=2,5$ or $10$ classes, where the data distribution is IID if $\sigma=10$, i.e., every device has all classes. The results are shown in Table 3, where we find that (i) FL training with non-IID data takes longer time to converge, and (ii) embracing both system and training efficiency utilities, WHALE-FL can remarkably improve FL training delay efficiency when applied to existing subnetwork methods under various data heterogeneity scenarios.
Conclusion